# Long Chain-of-Thought for Reasoning Large Language Models

### v3

@@ 3.2.3 Benchmarks
  • Commonsense Puzzle: Commonsense puzzle benchmarks, including LiveBench [676], BIG-Bench Hard [556] and ZebraLogic [359], assess models’ ability to reason about commonsense situations. The ARC [105] is often viewed as a challenging commonsense-based AGI test. JustLogic [75] further contributes to the evaluation of deductive reasoning and commonsense problem-solving. Moreover, Li et al. [307] introduce QuestBench, a benchmark designed to evaluate the ability of RLLMs to generate insightful and meaningful questions.
@@ 3.2.4 Process Evaluations
Recent progress in RLLMs underscores the need for specialized benchmarks to evaluate their deep reasoning abilities in Long CoT [301, 897]. Notably, Lin et al. [359] introduces ZebraLogic, a framework for assessing logical reasoning, especially in complex non-monotonic scenarios. Similarly, BigGSM [78] and GSM-Ranges [531] focus on perturbing numerical values to test logical and arithmetic reasoning in edge cases beyond the models’ training distribution. ROSCOE [166], ReCEval [481], and DiVeRSe [343] are designed to assess each step in the deep reasoning process during Long CoT tasks.
@@
Reflection benchmarks measure RLLMs’ ability to identify, reflect upon, and correct errors in Long CoT reasoning. These benchmarks fall into two categories: feedback and refinement. (1) Feedback Benchmark: These benchmarks assess the ability of LLMs to detect errors and respond to feedback for improvement. For example, Lambert et al. [294] introduces RewardBench to evaluate RLLMs’ reward capabilities. This framework is extended by Yasunaga et al. [760] and Zhang et al. [817] to include multimodal and code contexts, respectively. Benchmarks such as ProcessBench [871], PRMBench [538], MR-Ben [813], and DeltaBench [196] focus on error detection and correction across various tasks at the step level. Additionally, ReaLMistake [266] and JudgeBench [560] address more real-world error evaluation. (2) Refinement Benchmark: These benchmarks focus on error correction in complex tasks. CriticBench å[363] assesses critique-correction capabilities, while ErrorRadar [729] specializes in multimodal error detection, particularly in mathematics. FinerReason [61] introduces a commonsense puzzle for broader feedback and refinement evaluations. Medec [1] adapts error correction to healthcare, addressing medical issues.
@@ 3.2.5 Advanced Evaluation
  • Complex Science: This area involves integrating scientific texts with related diagrams or experimental data. Benchmarks like ScienceQA [399] and M3CoT-Science [79] evaluate how well models combine science information with Long CoT reasoning across various scientific domains [765]. Further, Guo et al. [181] propose MolPuzzle for the evaluation of molecular structure elucidation.
@@
Recent advancements in AI have significantly advanced scientific research [891, 653, 168], with platforms like SciWorld [636] improving the research process. Simultaneously, Pricope [483] and Chan et al. [58] introduce a machine-learning platform to evaluate the potential of RLLMs in automating experiments. Several studies also examine RLLMs’ ability to generate innovative research ideas. For instance, Si et al. [533] conduct evaluations with over 100 NLP researchers to assess RLLMs’ creativity, revealing notable limitations [326, 681, 576]. Additionally, Li et al. [349] introduce SolutionBench, a benchmark for assessing systems’ ability to generate feasible solutions for complex engineering problems.
@@ 4 Deep Reasoning for Long CoT
Deep reasoning capabilities primarily require profound depth and comprehensiveness in cognitive and reasoning processes. In the absence of such capabilities, RLLMs suffer significant performance declines [608, 659]. Current approaches to enhancing deep reasoning can be classified into two main categories: (1) Deep Reasoning Format (§ 4.1), which involves utilizing various reasoning execution formats, and (2) Deep Reasoning Learning (§ 4.2), which focuses on enabling the model to learn and strengthen its deep reasoning abilities.
@@ 4.1.2 Structured Language Deep Reasoning
Structured language deep reasoning encompasses various approaches designed to program [84, 371, 543, 470, 154, 673, 664] or symbolic language [480, 125, 360, 299, 739, 479, 32] format for enhanced deep reasoning. In this context, most studies focus on utilizing code to better enhance the mathematical reasoning capabilities [313, 91, 775, 73]. Xu et al. [711] propose a neural-symbol self-training framework guided by the environment, addressing both the scarcity of symbolic data and the limitations of symbolic processing in LLMs. Additionally, Liao et al. [353] present SKIntern, which refines symbolic RLLMs through curriculum learning and linear attenuation, enabling the internalization of symbolic knowledge with fewer examples, reducing computational costs, and accelerating inference. Furthermore, Ranaldi et al. [504] introduce QuaSAR, a CoT variant that directs LLMs to operate at higher abstraction levels through quasi-symbolic reasoning, thus improving natural language reasoning and providing more precise structural representations.
@@ 4.1.3 Latent Space Deep Reasoning
Latent space deep reasoning encompasses techniques designed to enhance the reasoning abilities of LLMs by leveraging operations within continuous latent spaces [541, 120, 510]. These approaches can be categorized into three main paradigms: (1) Reasoning Token-Driven Latent Space Deep Reasoning: Early work [646, 805] introduce the concept of “planning tokens” or “thought tokens” to guide reasoning within latent spaces. Further, Coconut [187] expands on this through the maintenance of multiple alternative reasoning paths, increasing both complexity and efficiency [849, 557]. At the extreme, Heima [528] condenses the entire Long CoT process into a single token, yielding substantial computational savings. (2) Reasoning Vector Driven Latent Space Deep Reasoning: Building on the previous paradigm, LTM [284] conceptualizes the layers of LLMs as “thought blocks” and introduces the concept of “thought vectors” for each layer. This approach allows for the scaling of test-time computations by implicitly performing reasoning within the latent space through recurrent depth. (3) Reasoning Manager Driven Latent Space Deep Reasoning: Inspired by these, Geiping et al. [159] and Saunshi et al. [516] propose a mechanism similar to a continuous reasoning manager, which iteratively governs a trained “recurrent block” as a recurrent “thought block”. This method integrates deeper model layers during reasoning, enhancing performance without needing specialized training data, and even outperforming larger RLLMs. Additionally, ITT [93] leverages the original transformer layer as a recurrent “thought block,” selecting key tokens via adaptive token routing and controlling reasoning depth with residual thinking connections, enabling more efficient processing of critical tokens.
@@ 4.2 Deep Reasoning Learning
Insufficient deep reasoning in RLLMs can significantly degrade performance [608, 659]. As a result, research has focused on improving reasoning through training. Supervised fine-tuning (SFT) [840] stabilizes model outputs by serving as a memory process [699], while reinforcement learning (RL) enables generalization and self-learning [179, 109]. Recent studies for deep reasoning learning have explored using SFT to imitate advanced reasoning in RLLMs and applying RL to enhance self-improvement in reasoning. As illustrated in Figure 6, this section outlines two key approaches to improve deep reasoning: (1) Deep Reasoning Imitation (§ 4.2.1), which involves learning reasoning from human-annotated or distilled data through SFT, and (2) Deep Reasoning Self-Learning (§ 4.2.2), where models improve reasoning through preference-based RL with implicit rewards. The performance of these methods is shown in Table 2.
@@ 4.2.1 Deep Reasoning Imitation
Deep reasoning in RLLMs can be effectively achieved by mimicking advanced reasoning systems, such as human reasoning [440, 53, 98, 325], advanced RLLMs [179, 51, 757, 297, 86], and Scaling-augmented RLLMs [332, 796, 473, 899, 35]. This approach enables the model to learn complex reasoning patterns and generalize across tasks [743]. Specifically, (1) Imitation from Human: Earlier, Cobbe et al. [113] first propose the deep reasoning imitation paradigm using human examples. ALT [440] improves RLLM reasoning by generating larger datasets of human-annotated logical templates, which fosters deeper reasoning. To enhance diversity, EIT [53] promotes simpler human-generated plans, while LLMs contribute more nuanced reasoning, facilitating collaboration between human input and AI. (2) Imitation from Advanced RLLMs: A body of work utilizes zero-shot prompting to guide large teacher RLLMs in generating reasoning rationale, which is then used to fine-tune smaller RLLMs, marking the beginning of deep reasoning imitation [199, 280, 769]. Additionally, AceMath [394] applies few-shot prompting to distill Long CoT samples from advanced LLMs, followed by multi-stage quality-guided SFT to enhance performance. Chen et al. [91] separate the data synthesis process into planning and reasoning stages, thereby improving reasoning quality. DART-Math [589] effectively distills complex queries requiring deeper reasoning during synthesis, advancing deep reasoning capabilities. Further, Ahmad et al. [6] propose OpenCodeReasoning, expanding this paradigm to the code scenarios. (3) Imitation from Scaling-augmented RLLMs: Earlier, Bansal et al. [30] enhance data quality by scaling the sampling size and length, boosting imitation performance. Yang et al. [734] and Zhao et al. [863] further improve data quality by scaling sampling and selecting samples through a reward model. Additionally, Li et al. [332] identify optimal deep reasoning paths through MCTS, advancing imitation effectiveness.
Recent studies [232, 435] show that distilling knowledge from advanced RLLM APIs like O1 [240] and R1 [179] significantly enhances the performance of smaller LLMs. This method, employing supervised fine-tuning, boosts model performance on complex mathematical reasoning tasks, sometimes surpassing the teacher models’ performance. Building on these findings, LIMO [766], S1 [441], and RedStar [715] argue that a large number of imitation samples is unnecessary. They demonstrate that even a minimal set of samples can activate deep reasoning capabilities in foundational LLMs. For practical applications, Turtel et al. [597] showcase how these techniques can predict future events beyond a model’s knowledge cutoff.
@@ 4.2.2 Deep Reasoning Self-Learning
While simple imitation can yield strong performance, current models still rely heavily on human annotations or outputs from more advanced models for both imitation and distillation [396]. To address this limitation, recent research has focused on enabling more advanced reasoning through techniques like self-play and self-learning [750, 855, 331]. Specifically, self-learning methods can be classified into two paradigms, differentiated by their sampling strategies:
@@ 5.1 Feedback
Feedback refers to the process of providing evaluations of both overall outputs and the processes that lead to them, with the goal of assessing their accuracy and quality [318, 320, 669, 173]. This process, also referred to as critique or verification, can be executed using either natural language or structured data formats, which serve as the foundation for tree-search methods [96]. Specifically, as shown in Figure 7, feedback can be categorized into three distinct types: (1) Overall Feedback (§ 5.1.1); (2) Process Feedback (§ 5.1.2); (3) Hybrid Feedback (§ 5.1.3).
@@ 5.1.1 Overall Feedback
Since many tasks cannot be directly evaluated using accuracy or other standard metrics, research has increasingly focused on Outcome Reward Models (ORM), which provide value-based rewards for more general and quantifiable feedback. In 2021, OpenAI [113] has proposed a “Gen-Verifier” paradigm, which uses a specialized ORM to evaluate the accuracy of generated rationales, showing significant progress in feedback capabilities [524]. Ji et al. [247] introduce a trained knowledge scorer to analyze hallucinations in the reasoning process, providing feedback to RLLMs and improving the accuracy of their outputs over time. Moreover, Generative Reward Models [834] use next-token prediction for overall feedback, which seamlessly integrates with instruction adjustments, leveraging test-time calculations to improve ORM feedback.
However, specifically trained ORMs are often costly and not sufficiently robust. Building on this, Self-Rewarding Language Models (SRLMs) [894] incorporate a self-consistency framework, optimizing feedback to improve model alignment and consistency. Yu et al. [785] introduce Critic-RM, combining RLLM-generated natural language criticism with corresponding feedback. This method filters high-quality feedback while jointly fine-tuning reward prediction and criticism generation, optimizing ORM performance.
@@
Although ORM has achieved significant improvements, its accuracy still falls short of 100%, preventing it from outperforming rule-based answer correction feedback [755, 185]. Previous studies, such as STaR [804], ReST [177], and ReFT Trung et al. [596], have demonstrated that feedback based on final answer rewards is more effective than both PRM and ORM in mathematical scenarios [153]. Furthermore, Guo et al. [179] and Xie et al. [701] introduce a multi-stage RL framework that incorporates rule-based rewards, significantly enhancing both output accuracy and length while mitigating reward hacking through simple yet robust rules [27], such as format validation and result verification. In coding scenarios where direct rule-based feedback is difficult, OpenCodeInterpreter [875], AceCoder [806], O1-Coder [854], and VerMCTS [49] address this challenge by implementing an automated test-case synthesis pipeline, deriving rewards based on program performance [445, 169, 882]. Additionally, Ma et al. [422] propose an automated approach to training a test case generator, which alleviates the scarcity of test cases and demonstrates that increasing the number of test cases correlates with improved reward quality. Moreover, Ma et al. [421] decompose problem-solving into structured subtasks: file localization, function localization, line localization, and code editing generation, and applies multi-viewed rule-based rewards.
@@
Research on feedback from RLLMs centers on detecting errors and biases through natural language feedback, also known as LLM-as-Judge, self-reflection or self-critique [214, 265, 26, 508, 434, 640, 795]. This method has led to significant improvements across various tasks, particularly in self-correction [674, 876, 161, 142, 853]. Huang et al. [223] contend that traditional LLMs struggle to generate effective feedback without external signals, requiring the development of RLLMs with enhanced feedback capabilities [515]. As a result, many studies leverage RLLMs’ error-identification strengths, often stemming from their pretraining phase, to improve feedback generation and correction [764, 34, 36].

### v4

@@ 3.2.3 Benchmarks
  • Commonsense Puzzle: Commonsense puzzle benchmarks, including LiveBench [818], BIG-Bench Hard [677] and ZebraLogic [427], assess models’ ability to reason about commonsense situations. The ARC [125] and DRE-Bench [912] is often viewed as a challenging commonsense-based AGI test. JustLogic [82] further contributes to the evaluation of deductive reasoning and commonsense problem-solving. Moreover, Li et al. [363] introduce QuestBench, a benchmark designed to evaluate the ability of RLLMs to generate insightful and meaningful questions.
@@ 3.2.4 Process Evaluations
Recent progress in RLLMs underscores the need for specialized benchmarks to evaluate their deep reasoning abilities in Long CoT [356, 1091]. Notably, Lin et al. [427] introduces ZebraLogic, a framework for assessing logical reasoning, especially in complex non-monotonic scenarios. Similarly, BigGSM [85] and GSM-Ranges [643] focus on perturbing numerical values to test logical and arithmetic reasoning in edge cases beyond the models’ training distribution. ROSCOE [198], ReCEval [580], DiVeRSe [404], HLV [67], and CoT-Kinetics [47] are designed to assess each step in the deep reasoning process during Long CoT tasks.
@@
Reflection benchmarks measure RLLMs’ ability to identify, reflect upon, and correct errors in Long CoT reasoning. These benchmarks fall into two categories: feedback and refinement. (1) Feedback Benchmark: These benchmarks assess the ability of LLMs to detect errors and respond to feedback for improvement. For example, Lambert et al. [348] introduces RewardBench to evaluate RLLMs’ reward capabilities. This framework is extended by Multimodal RewardBench[924], and CodeCriticBench [987] to include multimodal and code contexts, respectively. Benchmarks such as ProcessBench [1060], PRMBench [652], MR-Ben [983], and DeltaBench [234] focus on error detection and correction across various tasks at the step level. Additionally, ReaLMistake [318] and JudgeBench [681] address more real-world error evaluation. (2) Refinement Benchmark: These benchmarks focus on error correction in complex tasks. CriticBench [433] assesses critique-correction capabilities, while MLDebugging [270], and ErrorRadar [887] specializes in coding or multimodal reasoning error detection and refinement. FinerReason [68] introduces a commonsense puzzle for broader feedback and refinement evaluations. Medec [1] adapts error correction to healthcare, addressing medical issues.
@@ 3.2.5 Advanced Evaluation
  • Complex Science: This area involves integrating scientific texts with related diagrams or experimental data. Benchmarks like ScienceQA [484], M3CoT-Science [86], and ScienceBoard [670] evaluate how well models combine science information with Long CoT reasoning across various scientific domains [930]. Further, Guo et al. [215] propose MolPuzzle for the evaluation of molecular structure elucidation.
@@
Recent advancements in AI have significantly advanced scientific research [89, 1082, 787, 201], with platforms like SciWorld [768] improving the research process. Simultaneously, Pricope [582] and Chan et al. [63] introduce a machine-learning platform to evaluate the potential of RLLMs in automating experiments. Several studies also examine RLLMs’ ability to generate innovative research ideas. For instance, Si et al. [645] conduct evaluations with over 100 NLP researchers to assess RLLMs’ creativity, revealing notable limitations [384, 823, 698]. Additionally, Li et al. [413] introduce SolutionBench, a benchmark for assessing systems’ ability to generate feasible solutions for complex engineering problems.
@@ 4 Deep Reasoning for Long CoT
Deep reasoning capabilities primarily require profound depth and comprehensiveness in cognitive and reasoning processes. In the absence of such capabilities, RLLMs suffer significant performance declines [731, 793]. Current methods for enhancing deep reasoning can be categorized into two main approaches: (1) Deep Reasoning Format (§ 4.1), which involves utilizing various reasoning execution formats to maximize the reasoning step length $k$ within reasoning boundary $\mathcal{B}_{l}$ in Equation (2), by selecting the most suitable reasoning format; and (2) Deep Reasoning Learning (§ 4.2), which focuses on improving the model’s internal capabilities to enhance its deep reasoning abilities, thereby extending the reasoning boundary $\mathcal{B}_{l}$ in Equation (2) intrinsically.
@@ 4.1.2 Structured Language Deep Reasoning
Structured language deep reasoning encompasses various approaches designed to program [95, 441, 659, 565, 185, 813, 800, 1005] or symbolic language [579, 148, 428, 353, 898, 578, 34, 37, 767, 361] format for enhanced deep reasoning. In this context, most studies focus on utilizing code to better enhance the mathematical reasoning capabilities [369, 102, 941, 80]. Xu et al. [863] propose a neural-symbol self-training framework guided by the environment, addressing both the scarcity of symbolic data and the limitations of symbolic processing in LLMs. Additionally, Liao et al. [420] present SKIntern, which refines symbolic RLLMs through curriculum learning and linear attenuation, enabling the internalization of symbolic knowledge with fewer examples, reducing computational costs, and accelerating inference. Furthermore, Ranaldi et al. [608] introduce QuaSAR, a CoT variant that directs LLMs to operate at higher abstraction levels through quasi-symbolic reasoning, thus improving natural language reasoning and providing more precise structural representations.
@@ 4.1.3 Latent Space Deep Reasoning
Latent space deep reasoning encompasses techniques designed to enhance the reasoning abilities of LLMs by leveraging operations within continuous latent spaces [657, 143, 614, 305]. These approaches can be categorized into three main paradigms: (1) Reasoning Token-Driven Latent Space Deep Reasoning: Early work [780, 975] introduce the concept of “planning tokens” or “thought tokens” to guide reasoning within latent spaces [914, 970]. Further, Coconut [222] expands on this through the maintenance of multiple alternative reasoning paths, increasing both complexity and efficiency [1029, 678]. At the extreme, Heima [635] condenses the entire Long CoT process into a single token, yielding substantial computational savings. (2) Reasoning Vector Driven Latent Space Deep Reasoning: Building on the previous paradigm, LTM [337] conceptualizes the layers of LLMs as “thought blocks” and introduces the concept of “thought vectors” for each layer. This approach allows for the scaling of inference-time computations by implicitly performing reasoning within the latent space through recurrent depth. (3) Reasoning Manager Driven Latent Space Deep Reasoning: Inspired by these, Geiping et al. [191] and Saunshi et al. [620] propose a mechanism similar to a continuous reasoning manager, which iteratively governs a trained “recurrent block” as a recurrent “thought block” [488]. This method integrates deeper model layers during reasoning, enhancing performance without needing specialized training data, and even outperforming larger RLLMs. Additionally, ITT [104] leverages the original transformer layer as a recurrent “thought block”, selecting key tokens via adaptive token routing and controlling reasoning depth with residual thinking connections, enabling more efficient processing of critical tokens. Further, System-1.5 Reasoning [778] defines two dynamic shortcuts. The Model Depth Shortcut (DS) lets non-critical tokens exit early via lightweight adapter branches while routing critical tokens through deeper Transformer layers, thus supporting adaptive, vertical reasoning. The Step Shortcut (SS) reuses hidden states across decoding steps to bypass trivial iterations and enable horizontal reasoning in latent space.
@@ 4.2 Deep Reasoning Learning
Insufficient deep reasoning in RLLMs can significantly degrade performance [731, 793]. As a result, research has focused on improving reasoning through training. Supervised fine-tuning (SFT) [1018] stabilizes model outputs by serving as a memory process [849], while reinforcement learning (RL) enables generalization and self-learning [213, 131, 259, 864]. Recent studies for deep reasoning learning have explored using SFT to imitate advanced reasoning in RLLMs and applying RL to enhance self-improvement in reasoning. As illustrated in Figure 6, this section outlines two key approaches to improve deep reasoning: (1) Deep Reasoning Imitation (§ 4.2.1), which involves learning reasoning from human-annotated or distilled data through SFT, and (2) Deep Reasoning Self-Learning (§ 4.2.2), where models improve reasoning through preference-based RL with implicit rewards. The performance of these methods is shown in Table 2.
@@ 4.2.1 Deep Reasoning Imitation
Deep reasoning in RLLMs can be effectively achieved by mimicking advanced reasoning systems, such as human reasoning [533, 57, 110, 383], advanced RLLMs [213, 54, 921, 351, 97], and scaling-augmented RLLMs [390, 965, 570, 1094, 38]. This approach enables the model to learn complex reasoning patterns and generalize across tasks [902, 395]. Specifically, (1) Imitation from Human: Earlier, Cobbe et al. [135] first propose the deep reasoning imitation paradigm using human examples. ALT [533] improves RLLM reasoning by generating larger datasets of human-annotated logical templates, which fosters deeper reasoning [227]. To enhance diversity, EIT [57] promotes simpler human-generated plans, while LLMs contribute more nuanced reasoning, facilitating collaboration between human input and AI. (2) Imitation from Advanced RLLMs: A body of work utilizes zero-shot prompting to guide large teacher RLLMs in generating reasoning rationale, which is then used to fine-tune smaller RLLMs, marking the beginning of deep reasoning imitation [240, 333, 903, 498]. Additionally, AceMath [477] applies few-shot prompting to distill Long CoT samples from advanced LLMs, followed by multi-stage quality-guided SFT to enhance performance. Chen et al. [102] separate the data synthesis process into planning and reasoning stages, thereby improving reasoning quality. DART-Math [711] effectively distills complex queries requiring deeper reasoning during synthesis, advancing deep reasoning capabilities. Further, Ahmad et al. [6] propose OpenCodeReasoning, expanding this paradigm to the code scenarios. (3) Imitation from Scaling-augmented RLLMs: Earlier, Bansal et al. [32] enhance data quality by scaling the sampling size and length, boosting imitation performance [458, 967]. Yang et al. [892] and Zhao et al. [1048] further improve data quality by scaling sampling and selecting samples through sample feature or an additional reward model. Additionally, Li et al. [390] identify optimal deep reasoning paths through MCTS, advancing imitation effectiveness.
Recent studies [281, 526] show that distilling knowledge from advanced RLLM APIs like O1 [289] and R1 [213] significantly enhances the performance of smaller LLMs [403, 209]. This method, employing supervised fine-tuning, boosts model performance on complex mathematical reasoning tasks, sometimes surpassing the teacher models’ performance. Building on these findings, LIMO [931], S1 [534], and RedStar [868] argue that a large number of imitation samples is unnecessary. They demonstrate that even a minimal set of samples can activate deep reasoning capabilities in foundational LLMs. For practical applications, Turtel et al. [720] showcase how these techniques can predict future events beyond a model’s knowledge cutoff. Sun et al. [673], Yang et al. [893] and Zhao et al. [1051] further enhance deep reasoning imitation by selecting high-quality samples from large datasets, thereby improving the quality of the imitation data.
@@ 4.2.2 Deep Reasoning Self-Learning
While simple imitation can yield strong performance, current models still rely heavily on human annotations or outputs from more advanced models for both imitation and distillation [479]. To address this limitation, recent research has focused on enabling more advanced reasoning through techniques like self-play and self-learning [913, 1036, 389, 598]. Specifically, self-learning methods can be classified into two paradigms, differentiated by their sampling strategies:
@@ 5 Feasible Reflection for Long CoT
Feasible Reflection is a pivotal component of Long CoT reasoning, enabling LLMs to handle complex tasks through iterative feedback and refinement [386, 179]. Specifically, it comprises two primary stages: (1) Feedback (§ 5.1), which generates feedback signals $\mathcal{F}_{i}$ to correct node $n_{j}$ in Equation (5); and (2) Refinement (§ 5.2), which adjusts the subsequent node $n_{i+1}$ according to the feedback in Equation (LABEL:eq:refinement).
@@ 5.1 Feedback
Feedback refers to the process of providing evaluations of both overall outputs and the processes that lead to them, with the goal of assessing their accuracy and quality [374, 376, 808, 206, 829]. This process, also referred to as critique or verification, can be executed using either natural language or structured data formats, which serve as the foundation for tree-search methods [108]. Specifically, as shown in Figure 7, feedback can be categorized into three distinct types: (1) Overall Feedback (§ 5.1.1); (2) Process Feedback (§ 5.1.2); (3) Hybrid Feedback (§ 5.1.3).
@@ 5.1.1 Overall Feedback
Since many tasks cannot be directly evaluated using accuracy or other standard metrics, research has increasingly focused on Outcome Reward Models (ORM), which provide value-based rewards for more general and quantifiable feedback [1085, 949, 444]. In 2021, OpenAI [135] has proposed a “Gen-Verifier” paradigm, which uses a specialized ORM to evaluate the accuracy of generated rationales, showing significant progress in feedback capabilities [631]. Ji et al. [296] introduce a trained knowledge scorer to analyze hallucinations in the reasoning process, providing feedback to RLLMs and improving the accuracy of their outputs over time. Moreover, Generative Reward Models [1009] use next-token prediction for overall feedback, which seamlessly integrates with instruction adjustments, leveraging inference-time calculations to improve ORM feedback.
However, specifically trained ORMs are often costly and not sufficiently robust. Building on this, Self-Rewarding Language Models (SRLMs) [1087] incorporate a self-consistency framework, optimizing feedback to improve model alignment and consistency [1008]. Yu et al. [953] introduce Critic-RM, combining RLLM-generated natural language criticism with corresponding feedback. This method filters high-quality feedback while jointly fine-tuning reward prediction and criticism generation, optimizing ORM performance.
@@
Although ORM has achieved significant improvements, its accuracy still falls short of 100%, preventing it from outperforming rule-based answer correction feedback [919, 220, 1038]. Previous studies, such as STaR [974], ReST [211], and ReFT [718], have demonstrated that feedback based on final answer rewards is more effective than both PRM and ORM in mathematical scenarios [184]. Furthermore, Guo et al. [213] and Xie et al. [852] introduce a multi-stage RL framework that incorporates rule-based rewards, significantly enhancing both output accuracy and length while mitigating reward hacking through simple yet robust rules [29], such as format validation and result verification. In coding scenarios where direct rule-based feedback is difficult, OpenCodeInterpreter [1066], AceCoder [976], O1-Coder [1035], and VerMCTS [52] address this challenge by implementing an automated test-case synthesis pipeline, deriving rewards based on program performance [538, 202, 1073]. Additionally, Ma et al. [513] propose an automated approach to training a test case generator, which alleviates the scarcity of test cases and demonstrates that increasing the number of test cases correlates with improved reward quality. Moreover, Ma et al. [512] decompose problem-solving into structured coding subtasks: file localization, function localization, line localization, and code editing generation, and applies multi-viewed rule-based rewards.
@@
Research on feedback from RLLMs centers on detecting errors and biases through natural language feedback, also known as LLM-as-Judge, self-reflection or self-critique [257, 317, 28, 612, 525, 772, 964, 861]. This method has led to significant improvements across various tasks, particularly in self-correction [816, 1067, 193, 172, 1034]. Huang et al. [269] contend that traditional LLMs struggle to generate effective feedback without external signals, requiring the development of RLLMs with enhanced feedback capabilities [619, 378]. As a result, many studies leverage RLLMs’ error-identification strengths, often stemming from their pretraining phase, to improve feedback generation and correction [929, 36, 37, 265].